Weakly earnings on a certain import venture are approximately normally distributed with a known mean of $386 and unknown standard deviation. If the proportion of earnings over $419 is 25%, find the standard deviation. Answer only up to two digits after decimal.
Answer #1

We have given,

P[X>419]=0.25

Therefore,

P[Z>\frac{419-386}{\sigma}]=0.25

Now,

\frac{419-386}{\sigma}=0.67...............by using Z table or Excel =NORMSINV(1-0.25)

=>\sigma=49.25